OpenStudy (anonymous):

find the geometric mean of 125 and 5

OpenStudy (anonymous):

choices are 45 60 25 65

OpenStudy (anonymous):

and can you tell me how you got it please

OpenStudy (anonymous):

\[\sqrt{5 \times 125}=\sqrt{625}=25 \]

OpenStudy (anonymous):

thanks